OpenStudy (anonymous):

Slope Fields/ Separate Variables Find the equation of the curve, given the derivative and the indicated point on the curve. dy/dx=lnx/x

OpenStudy (anonymous):

point on graph is (1,-2)

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Same deal. Find the general antiderivatve of \(\dfrac{\ln(x)}{x}\).

OpenStudy (anonymous):

u mean integral?

OpenStudy (tkhunny):

Steve.

OpenStudy (anonymous):

so is it \[\ln \left| x \right| + C\]

OpenStudy (tkhunny):

Is that the right function? \(\dfrac{d}{dx}\ln|x| = \dfrac{1}{x}\). Whoops! That's not what we seek.

OpenStudy (anonymous):

so thats not the answer?

OpenStudy (tkhunny):

If you cannot find the derivative and get back to where you started, something went seriously wrong. \(\int \dfrac{\ln(x)}{x}\;dx = ??\)

OpenStudy (anonymous):

so would i use power rule?

OpenStudy (tkhunny):

You're just guessing, again. Ever hear of "The Substitution Method"?

OpenStudy (anonymous):

yes :)

OpenStudy (tkhunny):

Try u = ln(x) and see where it leads you.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

what would be the derivative of \[\frac{ u }{ \ln }\] so i can get du?

OpenStudy (tkhunny):

Never, EVER write just "ln" again. I would be tempted to fail you for the entire semester for that. "ln" is a function. It REQUIRES and argument. I am a little concerned. If you are on simple differential equations, you should have a firm grasp on the logarithm function and on integration by substitution. u = ln(x) du = (1/x)dx If you don't know this, you need to go back and find the section where you can re-learn it.

OpenStudy (anonymous):

so how did u get du

OpenStudy (anonymous):

ok i know what u is but how would i get x?

OpenStudy (tkhunny):

\(\int \dfrac{\ln(x)}{x}\;dx\) Given \(u = \ln(x)\;and\;du = \dfrac{dx}{x}\) This leaves \(\int u\;du\;=\;\dfrac{1}{2}u^{2} + C\) Finally, \(y = \dfrac{1}{2}(\ln(x))^{2} + C\) You MUST have seen this, before. Step up your game or you WILL fail this class.